OpenStudy (anonymous):

81^-3/4. Help, por favor!

OpenStudy (anonymous):

take fourth root, then cube, then invert.

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt[4]{81}^{3} }\]

OpenStudy (anonymous):

I got 1/27

OpenStudy (anonymous):

1/27 is right answer

OpenStudy (anonymous):

Awesome! Thank you.

OpenStudy (anonymous):

yes. the 4th root of 81 is 3. 3*3*3*3 = 81

OpenStudy (anonymous):

then 3^3=27

OpenStudy (anonymous):

so yes, 1/27 is your answer

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

:)